# Uncertainty Quantification for Evaluating Machine Translation Bias

### v1

@@ 4 Experimental setup
To test our proposed bias metrics, an MT dataset containing information about gender ambiguity and stereotypes in the source sentences is required. We use WinoMT (Stanovsky et al., 2019), which includes annotations on minimal pairs of 1,584 sentences with masculine, feminine or neutral pronouns referring back to stereotypical or anti-stereotypical gender roles. An example of three sentences from the dataset can be seen in Table 1. The gender of the focus noun ‘mechanic’ is unambiguous in the first two sentences based on the contextual information (Pronoun M & F), but remains ambiguous in the third on account of the neutral (N) pronoun. The Stereotype (column 3) that mechanics are more often men than women either contradicts (row 1) or aligns with (row 2) the disambiguating context (column 2).
@@
This dataset also contains other linguistic phenomena that were not explicitly annotated in its released version. Thus, we automatically annotate additional linguistic bias cues, namely subject, recency, implicit causality, and person names, using syntactic parses with Spacy2. We release the additional annotations to the public for reproducibility.
For the Subject bias, following the literature on human biases for coherence (Nieuwland and Van Berkum, 2006), we hypothesise that models may assume coreference between the subject of the main clause (often the focus noun) and the subject of the complement clause (often the pronoun). In the example in Table 1the mechanic’ is the subject, therefore the Subject bias primes an interpretation in which the gender of the subject aligns with that of the subsequent pronoun (M in row 1, F in row 2). Furthermore, person names have been shown to have a strong effect on pronoun resolution (Saunders and Olsen, 2023). To assess the impact of person names on gender translation, we augment the dataset with common feminine and masculine names matching the gender of the pronoun, selected for their cross-linguistic familiarity (see Appendix B). As an example, when translating into French, the second sentence from Table 1 would read “The mechanic Anne called to inform someone that she had completed the repair."
@@
Target Languages
We selected target languages that express gender through morphological markers on nouns and adjectives, and sometimes verbs, namely Spanish, French, Ukrainian, and Russian. The target languages vary in their representation within NLP research. A relevant idiosyncrasy of the Russian language is that for some nouns describing professions, even if a feminine version exists, it may be considered derogatory in use (Komova, 2024). For example, ‘врач’ is the masculine term for ‘doctor’, and the alternative feminine ‘врачиха’ is considered rude, which leads to ‘врач’ being used even when the doctor is known to be a woman. This results in constructions where masculine nouns are paired with feminine verb forms, or masculine markers are used throughout the sentence. In order to account for this, we also include a classification of professions which adhere to such (lack of) gender marking into WinoMT, using data from Komova (2024).
Human Translations
WinoMT does not contain target translations, thus the accuracy of MT models cannot be directly evaluated. To overcome this, we hired professional translators to translate a set of 100 WinoMT sentences into French, Spanish, Ukrainian and Russian. Each sentence is translated twice, with the focus noun in feminine and masculine variants respectively. They also annotate the translations as Correct or Incorrect with regard to the gender translation in the given context. For instance, when translating the English sentence “The farmer bought a book from the writer and paid her" into French, where ‘writer’ is the focus noun, the feminine ‘l’auteure’ should be marked as Correct, while ‘l’auteur’ would be Incorrect. In ambiguous cases, i.e. if the pronoun in the above sentence was ‘they’, both gender translations would be Correct. Appendix C provides the translation guidelines and details, and Appendix D discusses the quality of human annotations. We release the translations and correctness annotations to the public to enable further research.
Models
We experiment with two commonly used translation models, namely Opus-MT (Tiedemann and Thottingal, 2020) and m2m100 (Fan et al., 2021). Opus-MT models are NMT models trained on freely available parallel corpora. m2m100 is a many-to-many multilingual translation model, which directly translates between any pair of 100 languages. To examine how effective debiasing is regarding the three desiderata stated in Section 3, we apply the hard-debiasing method from Iluz et al. (2024) on the Opus-MT models, which have been shown to lower bias scores on the WinoMT dataset (Stanovsky et al., 2019) while maintaining translation quality. The hard-debiasing method neutralises the biased words in the representation space, so that neutral words are not associated with a specific gender (Bolukbasi et al., 2016). We adopt the most effective debiasing approach from Iluz et al. (2024), which applies debiasing to one-token profession words on the encoder side. See Appendix E for the performance of all models.
5 Research questions
Does Semantic Uncertainty Capture Gender Bias?
To validate the application of UQ metrics for bias evaluation, we compare their scores with the established gender accuracy metric. Gender accuracy uses the morphological parsers described in Section 3 to determine the focus noun gender in translations. As it relies on gold-standard references, it is applicable only to unambiguous items and unsuitable for cases with multiple valid gender realisations. We therefore limit this experiment to unambiguous items. We rank all models according to their $\Delta I$ scores and compare this ranking to that based on gender accuracy using Kendall’s $\tau$ and Pearson’s $r$. In order to establish whether a sampling-based metric is necessary, we also test a simple $\Delta LogProb$ value, which compares the Log Probabilities assigned to correct and incorrect instances, as an alternative to $\Delta I$.
What Biases do Models Exhibit via Uncertainty?
@@ 5 Research questions
To evaluate model bias, we assess how bias cues influence the diversity of gender markers in translations. Specifically, we perform an analysis of variance (anova) to examine the effect of bias cues (independent variables) on normalised entropy measures $-\mathcal{H}$ (s3e), $-\mathcal{H}$ (se), and $-\mathcal{H}$ (ge) (dependent variables), using T-tests for significance. This analysis incorporates linguistic biases not previously explored in WinoMT.
What does Semantic Entropy Reveal about Bias with Ambiguity?
6 Results
Lang. Model Gender Acc ΔLog $\Delta\text{Log prob}$ $\Delta I$ (s3e) comet ES Opus-MT 67.95 0.00 -0.10 84.90 deb-Opus-MT 68.13 0.00 -0.13 84.86 m2m100 70.77 0.00 -0.13 72.05 FR Opus-MT 64.27 0.01 -0.04 83.56 deb-Opus-MT 64.79 0.01 -0.08 83.55 m2m100 61.66 0.01 -0.07 73.06 UK Opus-MT 45.34 0.00 -0.03 70.79 deb-Opus-MT 46.12 0.00 -0.03 70.79 m2m100 47.76 0.00 -0.02 52.85 RU Opus-MT 48.57 0.00 0.00 79.37 deb-Opus-MT 48.42 0.00 -0.03 79.36 m2m100 48.49 0.00 -0.03 58.62
Table 2: Gender Accuracy, $\Delta$ Log Probability, and $\Delta I$ (s3e) on Unambiguous instances, comet scores on WMT test sets (see Appendix E for details).
Model Rankings According to Semantic Surprisal and Gender Accuracy Correlate.
The results of the first experiment are presented in Table 2. We find that while ΔLog $\Delta\text{Log prob}$ does not correlate with the gender accuracy ranking, $\Delta I$ (s3e) shows a statistically significant negative correlation with gender accuracy (Kendall’s $\tau=-0.58$, Spearman’s $\rho=-0.78$; see Appendix K). We attribute the effectiveness of this metric in distinguishing Correct from Incorrect gender translations to its ability to capture gender information beyond noun morphology, including verb inflections and agreement, through its embedding representations. The flexibility of s3e enables it to encode nuances that ge does not. For example, when translating a sentence with a feminine pronoun into Russian, Opus-MT generates sentences with a masculine noun and a verb that is either feminine or masculine (e.g., “Перевозчик поблагодарила (fem) / поблагодарил (masc)" “The courier thanked"). This is reflected in a higher $\mathcal{H}$ score by s3e (0.65) than ge (0.00), as the variation in verb inflections is only captured by s3e.
The strong negative correlation between $\Delta I$ (s3e) and gender accuracy thus validates the core component of our proposed metric for evaluating bias in machine translation.
In addition, the ranking of models by their overall performance, as indicated by comet scores, is partially aligned with the rankings based on Gender Accuracy and $\Delta I$ for unambiguous instances. This suggests that in these instances, better performing models tend to be less biased (all rankings are listed in Appendix J).

### v2

@@ 4 Experimental setup
To test our proposed bias metrics, an MT dataset containing information about gender ambiguity and stereotypes in the source sentences is required. We use WinoMT (Stanovsky et al., 2019), which includes annotations on minimal pairs of 1,584 sentences with masculine, feminine or neutral pronouns referring back to stereotypical or anti-stereotypical gender roles. An example of three sentences from the dataset can be seen in Table 1. The gender of the focus noun ‘mechanic’ is unambiguous in the first two sentences based on the contextual information (Pronoun M & F), but remains ambiguous in the third on account of the neutral (M/F) pronoun. The Stereotype (column 3) that mechanics are more often men than women either contradicts (row 1) or aligns with (row 2) the disambiguating context (column 2). [^3] Please see Appendix A for a broader overview of the biases present in the data.
Sentence Pronoun Stereotype
The mechanic called to inform someone that he had completed the repair. M M
The mechanic called to inform someone that she had completed the repair. F M
The mechanic called to inform someone that they had completed the repair. M/F M
Table 1: WinoMT Stanovsky et al. (2019) examples.
In addition, we experiment with the contemporaneously proposes dataset mGeNTe. The dataset consists of 1,500 sentences, with half of them containing unresolvably ambiguous gender instances, and half where the gender can be determined from the context. While this dataset does not explicitly mark stereotypical gender roles, it provides two alternative translations for each sentence, one with explicit gender marking, and one which avoids specifying the gender in the target language. Table 2 shows examples of both.
Sentence Gender Translation M/F Translation G
My second comment is that the previous speakers complained that the European Union pays, but the US is listened to. M/F In secondo luogo, le persone che mi hanno preceduto negli interventi hanno lamentato il fatto che l’Unione europea paga, mentre l’America si fa sentire. In secondo luogo, gli oratori che mi hanno preceduto hanno lamentato il fatto che l’Unione europea paga, mentre l’America si fa sentire.
The former Commissioner, Mr MacSharry, justified this criminal silence by claiming that the beef sector must not be put at risk. M MacSharry, in passato già esponente della Commissione, giustificava in tal modo questo silenzio colpevole, sostenendo che non si doveva mettere a repentaglio il settore bovino. L’ex Commissario MacSharry giustificava in tal modo questo silenzio colpevole, sostenendo che non si doveva mettere a repentaglio il settore bovino.
Table 2: mGeNTe Savoldi et al. (2025) examples.
The WinoMT dataset has annotations for the focus noun in the input sentence which is ambiguous with regard to gender, whereas in the mGeNTe the target sentences are annotated for gender on the relavant words.
Target Languages
We selected target languages that express gender through morphological markers on nouns and adjectives, and sometimes verbs, namely Spanish, French, Ukrainian, Russian, Italian, German and Greek. The target languages vary in their representation within NLP research. A relevant idiosyncrasy of the Russian language is that for some nouns describing professions, even if a feminine version exists, it may be considered derogatory in use (Komova, 2024). For example, ‘врач’ is the masculine term for ‘doctor’, and the alternative feminine ‘врачиха’ is considered rude, which leads to ‘врач’ being used even when the doctor is known to be a woman. This results in constructions where masculine nouns are paired with feminine verb forms, or masculine markers are used throughout the sentence. In order to account for this, we also incorporate a list of professions characterized by (lack of) gender marking in WinoMT, based on the categorisation provided by Komova (2024).
Human Translations
WinoMT does not contain target translations, thus the accuracy of MT models cannot be directly evaluated. To overcome this, we hired professional translators to translate a set of 100 WinoMT sentences into French, Spanish, Ukrainian and Russian. Each sentence is translated twice, with the focus noun in feminine and masculine variants respectively. They also annotate the translations as Correct or Incorrect with regard to the gender translation in the given context. For instance, when translating the English sentence “The farmer bought a book from the writer and paid her" into French, where ‘writer’ is the focus noun, the feminine ‘l’auteure’ should be marked as Correct, while ‘l’auteur’ would be Incorrect. In ambiguous cases, i.e. if the pronoun in the above sentence was ‘they’, both gender translations would be Correct. Appendix D provides the translation guidelines and details, and Appendix E discusses the quality of human annotations. We release the translations and correctness annotations to the public to enable further research.
Models
We experiment with two commonly used translation models, namely Opus-MT (Tiedemann and Thottingal, 2020) and m2m100 (Fan et al., 2021), as well as a multilingual LLM model Tower+ (Rei et al., 2025). Opus-MT models are NMT models trained on freely available parallel corpora. m2m100 is a many-to-many multilingual translation model, which directly translates between any pair of 100 languages. Tower+ is an open-weight model which not only exhibits high performance in MT, but also holds strong general capabilities. To examine how effective debiasing is regarding the expectations stated in Section 3, we apply the hard-debiasing method from Iluz et al. (2024) on the Opus-MT models, which have been shown to lower bias scores on the WinoMT dataset (Stanovsky et al., 2019) while maintaining translation quality. The hard-debiasing method neutralises the biased words in the representation space, so that neutral words are not associated with a specific gender (Bolukbasi et al., 2016). We adopt the most effective debiasing approach from Iluz et al. (2024), which applies debiasing to one-token profession words on the encoder side. Debiasing methods have been shown to have varying effects depending on the evaluation method (Mendelson and Belinkov, 2021; Gonen and Goldberg, 2019), which we also expect to see in our study. See Appendix F for the performance of all models.
5 Research questions
RQ1: Does Semantic Uncertainty Capture Gender Bias?
To validate the application of semantic UQ metrics for bias evaluation, we compare their scores with the established gender accuracy metric. Gender accuracy uses the morphological parsers described in Section 3 to determine the focus noun gender in translations. As it relies on gold-standard references, it is applicable only to unambiguous items and unsuitable for cases with multiple valid gender realisations. We therefore limit this experiment to unambiguous items. We rank all models according to their $\Delta I$ scores and compare this ranking to that based on gender accuracy using Kendall’s $\tau$ and Pearson’s $r$.
RQ2: What does Semantic Entropy Reveal about Bias with Ambiguity?
6 Results
RQ1: Model Rankings According to Semantic Surprisal and Gender Accuracy Correlate.
Lang. Model Gender Acc $\Delta\text{Log prob}$ $\Delta I$ (s3e) comet ES Opus-MT 67.95 0.00 -0.10 84.90 deb-Opus-MT 68.13 0.00 -0.13 84.86 m2m100 70.77 0.00 -0.13 84.53 Tower+9B 86.91 0.00 -0.30 89.10 FR Opus-MT 64.27 0.01 -0.04 83.56 deb-Opus-MT 64.79 0.01 -0.08 83.55 m2m100 61.66 0.01 -0.07 83.88 Tower+9B 75.77 0.02 -0.27 89.76 UK Opus-MT 45.34 0.00 -0.03 70.79 deb-Opus-MT 46.12 0.00 -0.03 70.79 m2m100 47.76 0.00 -0.02 76.52 Tower+9B 57.58 0.01 -0.12 92.26 RU Opus-MT 48.57 0.00 0.00 79.37 deb-Opus-MT 48.42 0.00 -0.03 79.36 m2m100 48.49 0.00 -0.03 80.49 Tower+9B 51.44 0.02 -0.09 90.92
Table 3: Gender Accuracy, $\Delta$ Log Probability, and $\Delta I$ (s3e) on Unambiguous instances of WinoMT, comet scores on WMT test sets (see Appendix F for details).
The results of the first experiment on WinoMT are presented in Table 3. We find that while $\Delta\text{Log prob}$ does not correlate with the gender accuracy ranking, $\Delta I$ (s3e) shows a statistically significant negative correlation with gender accuracy (Spearman’s $\rho=-0.84$; see Appendix M). We attribute the effectiveness of this metric in distinguishing Correct from Incorrect gender translations to its ability to capture gender information beyond noun morphology, including verb inflections and agreement, through its embedding representations. The flexibility of s3e enables it to encode nuances that ge does not. For example, when translating a sentence with a feminine pronoun into Russian, Opus-MT generates sentences with a masculine noun and a verb that is either feminine or masculine (e.g., “Перевозчик поблагодарила (fem) / поблагодарил (masc)" “The courier thanked"). This is reflected in a higher $\mathcal{H}$ score assigned to this sentence by s3e: $\mathcal{H}$ = 0.65, than ge: $\mathcal{H}$ = 0.00, as the variation in verb inflections is only captured by s3e.
The very strong negative correlation between $\Delta I$ (s3e) and gender accuracy thus validates the core component of our proposed metric for evaluating bias in machine translation. Our findings indicate that we can rank models according to their bias in unambiguous settings with our proposed metric as well as it can be done with gender accuracy, but without the need for explicit gender annotations.
We do not find a correlation between the surprisal scores and gender accuracy in the mGeNTe dataset (please see Appendix K for the scores). There are a number of factors that explain this. First, in WinoMT the Incorrect translations are explicitly the wrong gender, while in mGeNTe the Incorrect translations contain neutral expressions which do not match the explicit gender markers in the remainder of the same sentence (please see Translation N in the top example in Table 2). This is not strictly grammatically incorrect, only pragmatically less preferred, which is expected to elicit less surprisal from the model. Second, the mGeNTe dataset is not designed to include bias via stereotypes, as the gendered content of the sentence neither strongly aligns with nor contradicts the gender marking. This leads to smaller differences between the predicted likelihood of different gender expressions, with the highest absolute value of $\Delta H$ reaching 0.25, compared to 0.74 on WinoMT.
RQ2: Translation Accuracy Affects the Bias–Entropy Relationship Differently Across Levels of Analysis.
Having validated the s3e metric in the first experiment, we investigate the results of $\Delta\mathcal{H}$ (s3e) as a bias metric. Table 4 illustrates that all models exhibit the desired negative $\Delta\mathcal{H}$ for some languages in the WinoMT dataset. Tower+9B, which performs best in translation accuracy, also exhibits the desired unbiased behavior with consistently negative $\Delta\mathcal{H}$, suggesting that better models are less biased. Surprisingly, when it comes to ambiguous instances, contrary to the unambiguous cases, the languages where the models perform better on translation accuracy, namely Spanish and French, are not generally less gender biased under ambiguity (model rankings according to all metrics are available in Appendix L). Nonetheless, we observe the expected effect of debiasing on the Spanish and French translations (lower $\Delta\mathcal{H}$ for deb-Opus-MT).
Lang. Model Unamb Amb $\Delta\mathcal{H}$
ES Opus-MT 1.23 1.12 0.09
deb-Opus-MT 0.97 0.89 0.08
m2m100 1.79 1.45 0.19
Tower+9B 1.67 1.82 -0.15
FR Opus-MT 1.79 1.43 0.20
deb-Opus-MT 1.21 1.08 0.11
m2m100 3.22 2.78 0.14
Tower+9B 2.72 3.46 -0.74
UK Opus-MT 1.96 2.16 -0.10
deb-Opus-MT 1.98 2.15 -0.09
m2m100 2.05 2.28 -0.11
Tower+9B 4.12 4.85 -0.73
RU Opus-MT 1.56 1.68 -0.08
deb-Opus-MT 1.05 0.97 0.08
m2m100 1.83 2.29 -0.25
Tower+9B 3.15 3.63 -0.48
Table 4: Unambiguous and Ambiguous $\mathcal{H}$ (s3e) on WinoMT. Statistically significant correlations between $\Delta\mathcal{H}$ and comet scores on the instance level are in bold.
The results on the mGeNTe dataset corroborate the finding that the best performing model (Tower+9B) exhibits the least amount of bias (see Table 17 in Appendix K). The effects are smaller compared to the WinoMT dataset, and debiasing does not show an effect on mGeNTe, which may be explained by the fact that the dataset is not explicitly stereotypical and less bias can be observed.
Figure 2: Violin plots of $\mathcal{H}$ (s3e) on ambiguous and unambiguous WinoMT instances. Binned by low (B1), medium (B2) and high (B3) comet scores against human translations.
In Figure 2, results on WinoMT are grouped by comet score bins for a more fine-grained analysis at the instance level. For the models with negative $\Delta\mathcal{H}$ (s3e) scores, $\Delta\mathcal{H}$ is typically most pronounced for the highest-accuracy translations (e.g. ambiguous scores for Tower+9B-UK in Bin 3 are substantially higher than B1 and B2). Although debiasing does not reduce the overall $\Delta\mathcal{H}$ score for Ukrainian (see Table 4), it results in the largest improvement in the highest-quality translations: in bin B3, ambiguous $\mathcal{H}$ scores for deb-Opus-MT-UK are notably higher than those of the original model. This improvement is further supported by a substantial 8.41% drop in masculine focus noun inflections for Ukrainian, compared to 0.88–2.49% for other languages. We hypothesise that this is due to the limited training data in Ukrainian, which may lead to a less stable model that performs worse overall but is more responsive to debiasing in high-quality outputs. The relationship between translation accuracy and bias under ambiguity appears to differ depending on the level of analysis: between models, higher accuracy does not imply lower bias on ambiguous instances, whereas within models, higher-accuracy instances tend to show lower bias.
Qualitative Analysis
A qualitative analysis of the example in Table 1, presented with corresponding $\mathcal{H}$ values in Table 5, corroborates the quantitative findings in Figure 2. When translating the sentence in row 1, across all four target languages used in WinoMT, Opus-MT models consistently produce only the masculine variants of the focus noun (‘El mecánico’, ‘Le mécanicien’, ‘Механик’ and ‘Механiк’). In the anti-stereotypical case (row 2), all languages except Russian include both masculine and feminine forms (‘La mecánica’, ‘La mécanicienne’ and ‘Механiка’), indicating that these models are sensitive to the masculine stereotype even when the referent in the context is clearly feminine. For Russian, the models fail to generate any feminine constructions, even when the context is unambiguously feminine. This difference is evident in the $\mathcal{H}$ (s3e) scores, which are higher in row 2 than row 1 for the first three languages. Moreover, in the ambiguous case (row 3), all Opus-MT models produce only masculine nouns, regardless of language. Consequently, $\mathcal{H}$ (s3e) scores are generally higher for the unambiguous cases (mean of rows 1 and 2) than for the ambiguous case (row 3), except for Russian, where $\mathcal{H}$ remains low across all conditions. These observations are consistent with expectations for biased models: they default to stereotypical gender realisations when the pronoun is ambiguous and sometimes even when the context clearly suggests an anti-stereotypical interpretation.
7 Conclusion
Limitations
Ethics Statement
Appendix A Effects of Different Biases
@@
This dataset also contains other linguistic phenomena that were not explicitly annotated in its released version. Thus, we automatically annotate additional linguistic bias cues, namely subject, recency, implicit causality, and person names, using syntactic parses with Spacy(https://spacy.io/). We release the additional annotations to the public for reproducibility.
For the Subject bias, following the literature on human biases for coherence (Nieuwland and Van Berkum, 2006), we hypothesise that models may assume coreference between the subject of the main clause (often the focus noun) and the subject of the complement clause (often the pronoun). In the example in Table 6the mechanic’ is the subject, therefore the Subject bias primes an interpretation in which the gender of the subject aligns with that of the subsequent pronoun (M in row 1, F in row 2). Furthermore, person names have been shown to have a strong effect on pronoun resolution (Saunders and Olsen, 2023). To assess the impact of person names on gender translation, we augment the dataset with common feminine and masculine names matching the gender of the pronoun, selected for their cross-linguistic familiarity (see Appendix C). As an example, when translating into French, the second sentence from Table 6 would read “The mechanic Anne called to inform someone that she had completed the repair."
@@
To evaluate model bias, we assess how bias cues influence the diversity of gender markers in translations. Specifically, we perform an analysis of variance (anova) to examine the effect of bias cues (independent variables) on normalised entropy measures norm-$\mathcal{H}$ (s3e), norm-$\mathcal{H}$ (se), and norm-$\mathcal{H}$ (ge) (dependent variables), using T-tests for significance. This analysis incorporates linguistic biases not previously explored in WinoMT.
Our expectation is that the entropy of unbiased models should not be affected by the presence of bias cues. Thus, we define normalised entropy, which compares the $\mathcal{H}$ of a source sentence $x$ to the average entropy across its contrast set $\mathcal{G}_{x}$. $\mathcal{G}_{x}$ is a group of minimally different sentences that are identical to $x$ apart from the pronoun (e.g., ‘she’, ‘he’, ‘they’), including $x$ itself. The three sentences in Table 1 comprise a $\mathcal{G}_{x}$. Formally:
$\text{norm-}\mathcal{H}(x)=\frac{\mathcal{H}(\mathcal{Y}_{x})}{\frac{1}{|\mathcal{G}_{x}|}\sum_{x^{\prime}\in\mathcal{G}_{x}}\mathcal{H}(\mathcal{Y}_{x^{\prime}})},$ (6)
The results of the experiment presented in Table 7 show that most bias cues in the data have a significant effect on the variance of norm-$\mathcal{H}$ (s3e), indicating that the tested models exhibit various social and linguistic biases.444norm-$\mathcal{H}$ (s3e) shows the strongest sensitivity to bias cues compared to norm-$\mathcal{H}$ (se) and norm-$\mathcal{H}$ (ge). We also experiment with unnormalised $\mathcal{H}$ scores, the results of which are less comparable across metrics, bias types, and models. The full results are presented in Appendix G. All trends observed for norm-$\mathcal{H}$ (s3e) are also present in norm-$\mathcal{H}$ (se) and norm-$\mathcal{H}$ (ge), as well as for unnormalised $\mathcal{H}$. The results corroborate previous findings. The high absolute coefficient values in the Names column indicate that person names have an effect on gender translation even when a disambiguating pronoun is present. This is in line with Saunders and Olsen (2023), who have shown that both pronouns and names induce gender bias and are often not sufficient for full disambiguation. Secondly, the fact that some Russian nouns have a default masculine grammatical gender regardless of the context (Komova, 2024) is reflected in significant decrease in gender diversity for sentences containing such nouns (negative coefficients in the Default M columns indicating lower norm-$\mathcal{H}$). Thirdly, we observe that masculine biases generally reduce norm-$\mathcal{H}$ (negative coefficients in the M columns), while feminine biases tend to increase it (positive coefficients in the F columns). This suggests a general default toward masculine translations in models, with outputs becoming more similar under masculine biases and more varied under feminine ones. This finding aligns with Kuzucu et al. (2025), who show that model uncertainty is typically higher for minority groups. The results in the Ambiguity column in Table 7 indicate that norm-$\mathcal{H}$ increases for Ukrainian and Russian (positive coefficients), but not for Spanish and French, which mirrors the results from Section 6. Higher norm-$\mathcal{H}$ for Ambiguous items (or negative $\Delta\mathcal{H}$) is expected for an unbiased model.